OpenStudy (anonymous):

Solve the equation x − 6 = 11

OpenStudy (anonymous):

add 6 to both sides

OpenStudy (anonymous):

Add 6 to both the sides and you will have your answer..

OpenStudy (hba):

17 is the answer

OpenStudy (hba):

\[x=11+6\]

OpenStudy (anonymous):

x-6=11 x=11+6 x=17